# Constant search time algorithm via topological quantum walks

### v1

@@ Abstract
It is well-known that quantum algorithms such as Grover’s can provide a quadradic speed-up for unstructured search problems. By adding topological structure to a search problem, we show that it is possible to achieve a constant search-time quantum algorithm with a constant improvement of the search probability over classical search. Specifically, we study the spatial search algorithm implemented by a two-dimensional split-step quantum random walks that realize topologically nontrivial phases and show the asymptotic search behavior is constant with growing system size. Using analytical and numerical calculations, we determine the efficient search regions in the parameter space of the quantum walker. These regions correspond to pairs of trapped states formed near a lattice defect. By studying the spectral properties of the discrete time-evolution-operators, we show that these trapped states have large overlap with the initial state. This correspondence, which is analogous to localization by constructive interference of bound states, makes it possible to reach the best possible search-time asymptotic and produce a disorder-protected fast search in quantum random walks.
Quantum search algorithm, Grover’s search [1], is known as an algorithm that demonstrates quantum advantage over its classical analogue. Grover’s algorithm performs a search of marked element in an unstructured database with a given probability with asymptotic time being a square root of database element count. However, its implementation depends on an oracle operator - a way of distinguishing marking element that applies to a whole quantum state. The quantum random walks are known as an example of natural implementation of such an oracle operator [2, 3, 4, 5, 6]. The discrete-time quantum random walks are defined as a set of shift and spin flip operators applied at the whole system at each time step. The successful search event is defined as a time of maximum localization of quantum walker on a defect when starting from a uniform distribution over all system.
Although, different kinds of quantum walks demonstrate a Grover-type convergence of a spatial search in two and higher dimensions, the dependence on spin flip operator and position of defect site relative to remaining system is crucial and in many cases results in a decreased efficiency of search algorithm [5, 4, 7]. The latter fact also serves as a motivation for the search for sub-optimal yet more robust to disorder and element positioning algorithms, that would provide quantum advantage. In the present paper we study the two-dimensional split-step quantum random walks that are analogues of Chern topological insulators [8] in crystalline solids. The similar split-step QRW in 1D were studied in Refs.[9, 10] and a number of our conclusions made below might work also in those cases. By analyzing the parameter spaces of quantum walker and defect, we find that there exist sets of parameters that demonstrate efficient search of a given defect (see a search probability map in Fig.1). Below we show that these regions in parameter space can be categorized into two distinct classes based on their position inside the topological phase or on the phase separation lines. Notably, the search time for the parameters inside both classes of regions demonstrates saturation to a constant search time with growing system size. The saturation of search time happens at different system sizes, while before that the square root time growth behavior takes place.
@@
Figure 1: The success probability of search algorithm as a function of QRW parameters $\Theta_{1,2}$ (see Eq.(2)) for a defect with $\Theta_{1}^{def}=5\pi/8$, $\Theta_{2}^{def}=\pi/2$. High peaks are of order of magnitude larger than average localization probability and correspond to parameters for which quantum walker performs search efficiently. Inset: geometry of split-step QRW on a triangular lattice with three shift vectors ${\boldsymbol{v}}_{i}$ solid ones correspond to shift of spin up components, dashed $-{\boldsymbol{v}}_{i}$ for spin down.
The constant search time asymptotic was previously indentified only for electric Dirac QRW with non-local marking of defect in the form of Coulomb potential [11, 12]. The non-locality creates an important difference from our algorithm. Namely, here the constant search time is achieved with local marking, thus making it directly implementable in experimental memory setup.
@@
Below we analyze in detail the time evolution of states for a range of quantum walker parameters and study the overlap of initial state with the eigenstates of a unitary operator corresponding to a single time step of QRW. We demonstrate that the search speedup with constant time asymptotic is related to the appearance of trapped states of a specific structure near the defect: the two pairs of trapped states should simultaneously have a large matrix element of coordinate operator of the defect and the large overlap with initial uniformly distributed state. In the mathematical sense the trapped states are defined as a states that have compact support and zero density elsewhere. Their analogue in quantum mechanics are the strongly localized bound states on impurity in a system. We note that a number of mathematically rigorous studies have been made for spectral properties of quantum random walks and trapped states [5, 13, 9, 14, 15, 16] together with their dynamics [17, 18]. However, the possible efficiency of search algorithm based on exploiting the particular structure of QRW was not previously addressed. The strong dependence of the trapped states structure on defect and walker parameters also suggests a scenario of protecting quantum search from a random disorder by an appropriate fine-tuning of marking parameters for the searched element.
@@
The quantum random walk deployed for algorithmic speed-up in this work was firstly introduced in Ref.[8] as a way to simulate topological insulators. Its single time step
@@
is defined via three rotations of spin and subsequent translations:
$U(\Theta_{1},\Theta_{2})=T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right)T_{{\boldsymbol{v}}_{2}}R\left(\theta_{2}\right)T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right),$ (2)
@@
Figure 2: (a) A colormap representation of a maximum localization probability on a defect site over the evolution time $T=1000$ for a system of size $N=40\times 40$ starting from a fully uniform distribution. Black numbers indicate Chern numbers of the corresponding parameter regions and dashed lines correspond to phase separation lines. Cyan dot marks the defect parameters $(5\pi/8,\pi/2)$. (b) Search time measured as a position of first probability peak as a function of system size for walker parameters indicated by crosses ‘1’ and ‘2’ in panel (a). (c) Search probability dependence on system size. The two regions of different dependence are present where probability follows Grover-type scaling $P\sim 1/\log(N)$ until the time asymptotic changes to constant in (b) and then switches to system size scaling $P\sim 1/N$ but having few orders of magnitude larger value than average density in the system. (d) Comparison of performance with Grover’s algorithm and repeated classical search for system size $\sqrt{N}=512$. Walker $2$ shows advantage over classical.
In our setup the search algorithm goal is defined as a problem of finding a single node on a finite 2D lattice of size $\sqrt{N}\times\sqrt{N}$ unit cells starting from initial state $|i\rangle$. We apply the periodic boundary condition in order to avoid the appearance of edge states. The searched node is marked by parameters $\Theta_{1,2}^{def}$ that are different from surrounding $\Theta_{1,2}$. To estimate the possibility of given quantum walker to perform search efficiently, we firstly scan the entire parameter range $\Theta_{1,2}$ for a given defect values $\Theta_{1,2}^{def}$. The performance of search algorithm is measured by a search probability. It is defined as a peak $\max_{t\in[0,T]}P_{def}(t)$ in time-dependent density of the system state at the defect node,
@@
where $|d\rangle=\sum_{\sigma=\uparrow,\downarrow}|x=def,\sigma\rangle$ is the defect coordinate operator with both components of spin $\sigma$, and $|i\rangle$ is the initial state of the system defined as $|i\rangle=\frac{1}{\sqrt{2N}}\sum_{\sigma,x}|x,\sigma\rangle$ with summation over entire system. The plot of the maximum value of density $P_{def}(t)$ on defect over the interval of time $T=1000$ for test system size $N=40\times 40$ sites is shown in Fig.1 for $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$. In Fig.2(a) we show the colormap representation of the same search probability distribution on the top of phase diagram in parameter space (see also Supplement [21] for other defect parameters).
@@
The islands of efficient search in parameter space fill a small areas placed in two distinct types of regions: inside topological phases and on phase separation lines. We compare the search time and probability behavior with growing system size for the walker parameters marked by ‘1’ and ‘2’ in the Fig.2(a), that realize best search probability within its regions. Such results are shown in Fig.2(b). With the growing system size the behavior of search time defined via first probability peak demonstrates qualitative change - from $T\sim\sqrt{N}$ dependence to $T\sim const$ in both curves. As we noted previously, such speed-up was observed only for the Dirac QRW with particular type of Coulomb-like disorder potential [11, 12]. In Fig.2(c) the search probability dependence of walkers ‘1’ and ‘2’ on system size is compared with average density at non-defect nodes and Grover’s algorithm search probability. We find that the constant search time regime corresponds to the $P_{def}(t)=O(1/N)$ with a large numerical prefactor. This decrease of search probability with system size $N$ is faster than Grover’s one, $P_{def}^{Grover}=O(1/log\,N)$. The search probability dependence in the form $O(1/N)$ arises due to the trapped states having finite support. The system size grows beyond the trapped state support size corresponds to saturation of search time asymptotic to constant and clear change in probability dependence curve in Fig.2(c). To check the dependence of search probability on defect parameters for given walker we performed the same 2D parameter space calculation taking walker "$1"$ and checking $\Theta_{1,2}^{def}\in[-2\pi,2\pi]$. It is found that the search probability does not depend on $\Theta_{2}^{def}$ [21], thus the more informative is to present a dependence on $\Theta_{1}^{def}$ as shown in Fig.3(a).
The asymptotic scaling of $O(1/N)$ in search probability does not reach the optimal Grover scaling, but still provides constant improvement over the classical search algorithm. The probability of classical algorithm grows as $t/N$ with multiple trials, $t$ (one per time step). For this purpose, in Fig.2(d) we present a comparison of probability on searched node as a function of time for two walkers marked in Fig.2(a), Grover’s search for split-step QRW from Ref.[7] and a classical search. The walker ‘2’ demonstrates twice as large probability as classical search, while the walker ‘1’ follows the classical search probability for approximately 100 steps. In terms of detection via measurements, higher probability on a node means that the detector resolution used to identify position could be lower - in particular, for walker ‘1‘ the detection resolution could be $100$ times smaller that for classical probes of individual nodes. The Grover’s algorithm performance demonstrates even higher probability at the same time scale, which makes it advantageous in the non-disordered system. As we point out below, that advantage of Grover’s search might be lost in a disordered systems.
@@
To describe the origin of different speed-ups in this system and why the regions of efficient search represent a small fraction of parameter space, we analyze the spectrum and localization properties of eigenvectors of a unitary operator of system with one defect. The probability evolution at a defect site at discrete times $t$ can be rewritten through the unitary operator of QRW,
@@
$P_{def}(t)=\bigg{|}\langle d|\sum_{n,\lambda=\pm}e^{i\lambda E_{n}t}|n,\lambda\rangle\langle n,\lambda|i\rangle\bigg{|}^{2}$ (8)
@@
The saturation of search time to constant scaling after system size becomes large enough suggests that the properties of states that mainly contribute to the Eq.(8) do not change with system size. This feature directly corresponds to the definition of trapped states previously introduced in Refs. [5, 13] in the context of QRWs staying localized around its initial starting node over infinite evolution time. We investigate the contribution of different eigenstates of $\bar{U}$ to the evolution of probability, $P(t)$, in Fig.3(b) by comparing the product of overlaps $|\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$. The panels in Fig.3(b) correspond to different parameters of defect denoted by vertical dashed lines for walker `$1^{\prime}$ in Fig.3(a).
@@
Figure 3: (a) Comparison of maximum value of probability on defect for a walker ‘1’ from Fig.2 as a function of defect parameter $\Theta_{1}^{def}$ with the maximal value of squared of overlaps defined in Eq.(9) with $\mathcal{M}$ containing one and two states respectively. All quantities are normalized to scale (0,1) for the purpose of comparison. The summation over two trapped states gives very accurate approximation of search probability peak. (b) The squared product of overlaps plotted as a function of energy of each eigenstate of operator $\bar{U}$ for walker parameters ‘1’ and defect parameter $\Theta_{1}$ indicated on top of each panel. Trapped states differ from other states by significantly higher value of overlap.
@@
The scale of $x$-axis in each panel of Fig.3(b) shows that defect parameters with higher search probability also correspond to higher overlap product for a few states - in this case two pairs of particle-hole symmetric states. To further verify this observation, we compare the maximal value of overlap product
$W_{\mathcal{M}}=|\sum_{n\in\mathcal{M},\lambda=\partial}\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$ (9)
@@
of a single eigenstate of $\bar{U}$ with the normalized probability dependence $max[P_{def}(t)]$ as function of $\Theta_{1}^{def}$. The comparison is made by taking a single pair of states that maximize overlap with set of indices =$\mathcal{M}=\{n_{max}\}$, as well as two pairs of states =$\mathcal{M}=\{n_{max},n_{second\,max}\}$. The results presented in Fig.3(a) indicate that while =$\mathcal{M}=\{n_{max}\}$ selection overestimates the region of good search, the =$\mathcal{M}=\{n_{max},n_{second\,max}\}$ very precisely describes the high search probability peak. To further check this correspondence between the evolution of described by two pairs of states with maximal overlap and the max probability distribution, we plotted the same comparison for a defect $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$ along phase separation line and the line with walker `$1^{\prime}$ in Supplemental material. The correspondence for a test system size with $\sqrt{N}=40$ shows that efficient search for such split-step QRW is mostly determined by the appearance of two pairs of trapped states having large overlap with both defect and initial state.
@@
where we took into account the particle-hole symmetry relation $w_{-,j}^{\sigma}=(w_{+,j}^{\sigma})*$. It is important to emphasise that after calculating probability at a defect site, one finds that the two spin components exactly cancel oscillating parts of each other for a single value of $j$ due to the different phase of oscillations between $sin$ and $cos$ in equations (10),(11) above
@@
with $\text{sign}(\sigma)=\pm$. Thus, the contribution of each single pair of trapped states is constant and due to the initial condition of uniformly distributed state over entire system it could not describe the peak in probability.
Figure 4: The comparison of density evolution at searched defect node (blue solid line) and another distanced node without defect (gray dashed line) with time for a walker parameters ‘1’ and defect marked by $(5\pi/8,\pi/2)$. The red dash-dotted line corresponds to the evolution described by only two pairs of trapped states that maximize product of overlaps and given by Eq.(Constant search time algorithm via topological quantum walks). The evolution from trapped states approximates with very good precision the first few peaks in probability and the time of their occurence.
@@
The period of this oscillation is different from the fast oscillations of each trapped states and defined by the energy difference between the positive energy levels of two pairs of trapped states $(E_{1}-E_{2})$. The plot of evolution of probability on defect in Fig.4 shows that such simple expression captures very well the structure of oscillations, the time of first peak appearance and the max probability on the first period on a defect. In Fig.4 the results are presented for a walker ‘1’, and the same is shown for a walker ‘2’ in Supplemental material. An another way to check the agreement of predicted period of oscillations with the actual oscillation of probability on a defect node we used numerical Fourier transform. The agreement with period defined by $(E_{1}-E_{2})$ has errors below $2\%$ on the timescale of evolution $T=5000$ for system with $\sqrt{N}=40$ [22]. Note that such definition of period through the difference of eigenstates of operator is different from the common one used in Grover’s search in QRW [23, 7]. Specifically, the period definition we use here is between states from different particle-hole symmetric pairs rather that between states from single pair.

### v3

# Dynamical localization in 2D topological quantum random walks

@@ Abstract
We study the dynamical localization of discrete time evolution of topological split-step quantum random walk (QRW) on a single-site defect starting from a uniform distribution. Using analytical and numerical calculations, we determine the high localization probability regions in the parameter space of the quantum walker. These regions contain two or more pairs of trapped states, forming near a lattice defect. By investigating the spectral properties of the discrete-time evolution operators, we show that these trapped states have large overlap with the initial uniformly distributed state, thus offering a simple interpretation of the localization effect. As this localization scheme could be interpreted as a variation of spatial quantum search algorithm, we compare the localization probability and time with other types of two-dimensional quantum walks that do not have topological phases and realize localization time scaling similar to Grover’s algorithm. Finally we show that mechanism of localization we identified is robust against the influence of disorder.
I Introduction
The quantum random walks were introduced in Ref. [1] as a natural quantum version of classical random walks. Due to interference effects and the state superposition, the average travel path of walkers is much longer than their classical counterpart. The clear structure of possible physical systems that could realize the quantum walks [1] attracted much attention. In particular, it was shown that quantum walks are a possible way to implement Grover’s algorithm [2] in the case of absence of high-quality multi-qubit quantum computer [3, 4, 5, 6, 7, 8]. It was also shown that the so-called Grover’s oracle operator has a natural realization in quantum walks and is related to dynamical localization phenomena [3, 4, 5, 6, 7, 8]. In particular, the discrete-time quantum random walks are defined as a set of shift and coin-flip operators applied to the whole system at each time step [5, 9, 1, 3]. The successful search event is defined as a time of maximum localization of quantum walker on a defect when starting from a uniform distribution over the whole system. Going forward, we will refer to the the time of appearance of the first maximum in localization probability as search time.
@@ I Introduction
Figure 1: The localization probability of topological quantum walk as a function of spin rotation parameters $\Theta_{1,2}$ (see Eq. (2)) for a defect with $\Theta_{1}^{def}=5\pi/8$, $\Theta_{2}^{def}=\pi/2$. High peaks are of two orders of magnitude larger than average localization probability and correspond to parameters for which quantum walker performs search efficiently. Inset: geometry of split-step QRW on a triangular lattice with three shift vectors ${\boldsymbol{v}}_{i}$ solid ones correspond to shift of spin up components, dashed $-{\boldsymbol{v}}_{i}$ for spin down.
In the most common mechanism of localization, at least several eigenstates of QRW evolution operator create a constructive superposition on a defect at a certain time [6, 5]. Little is known about the precise values of optimal parameters of defect that maximize localization relative to the surrounding system. In addition, it was shown that certain positioning of defects could significantly decrease localization probability [10]. Here, we study the two-dimensional split-step quantum random walks analogous to Chern type topological insulators in crystalline solids [11]. For such a quantum walk, it could be expected that topology creates a new mechanism of localization, similar to in-gap states in solids. Following this motivation we explore how topology affect the localization mechanism. The topological split-step QRWs in 1D space were studied in Refs.‘[12, 9]. Here, we analyze the parameter spaces of the quantum walker and the defect. We find that there exist sets of parameters that deterministically demonstrate a high localization probability at a given defect (see a localization probability map in Fig. 1). In Sec. II we show that these regions in parameter space can be categorized into two distinct classes based on their position inside the topological phase or on the phase separation lines. Notably, we find the search time for the parameters inside both classes of regions demonstrates the saturation to a constant with growing system size. The saturation of search time happens at different system sizes, while before this saturation the square root time growth behavior takes place.
The localization after constant time, regardless of system size, was previously identified only for electric Dirac QRW with non-local marking of defect in the form of Coulomb potential [13, 14]. The non-locality of defects creates an important difference from our scenario: our protocol does not require a large number of nodes to form a defect structure. In the algorithmic terminology, in our work the constant search time is achieved with local marking comparing to non-local extended marking of Coulomb potential.
@@
In Sec. III we analyze the time evolution of QRW states for a range of parameters in detail, and study the overlap of initial state with the eigenstates of the unitary operator corresponding to a single time step of QRW. Particularly, we demonstrate that the search speedup with constant time asymptotic is related to the appearance of trapped states of a specific structure near the defect: the two pairs of trapped states should simultaneously have a large matrix element with coordinate operator of the defect, and have large overlap with uniformly distributed initial state. In mathematical terms the trapped states are defined as a states that have compact support and zero density elsewhere. Their analogue in condensed matter are the strongly localized bound states on impurity in a system. We note that a number of mathematically rigorous studies have been made for spectral properties of quantum random walks and trapped states [6, 15, 12, 16, 17, 18] together with their dynamics [19, 20]. However, the possible efficiency of search algorithm based on exploiting the particular structure of QRW was not previously addressed. The strong dependence of the trapped states structure on defect and walker parameters also suggests a scenario of protecting quantum search from a random disorder by an appropriate fine-tuning of marking parameters for the searched defect element. The corresponding study is presented in Sec. IV and the concluding remarks are given in Sec. V.
II The QRW model and localization phenomenon
@@
The quantum random walk analyzed in this work was first introduced in Ref.[11] as a way to simulate topological insulators. Its single time step
@@
is defined via three rotations of the spin and subsequent translations:
$U(\Theta_{1},\Theta_{2})=T_{{\boldsymbol{v}}_{3}}R\left(\theta_{1}\right)T_{{\boldsymbol{v}}_{2}}R\left(\theta_{2}\right)T_{{\boldsymbol{v}}_{1}}R\left(\theta_{1}\right),$ (2)
@@
Figure 2: (a) A color map representation of a maximum localization probability on a defect site over the evolution time $T=1000$ for a system of size $N=40\times 40$ starting from a fully uniform distribution. Black numbers indicate Chern numbers of the corresponding parameter regions and dashed lines correspond to phase separation lines. Cyan dot marks the defect parameters $(5\pi/8,\pi/2)$. (b) Search time measured as a position of first probability peak as a function of system size for walker parameters indicated by crosses ‘1’ and ‘2’ in panel (a). (c) Localization probability dependence on system size. The two regions of different dependence are present where probability follows ‘Grover-type’ scaling $P\sim 1/\log(N)$ as non-topological QRW from Ref.[10] until the time asymptotic changes to constant in (b) and then switches to system size scaling $P\sim 1/N$ but having few orders of magnitude larger value than average density in the system. (d) Comparison of time-dependent localization probability with non-topological QRW [10] and repeated classical search for system size $\sqrt{N}=512$. Walker $2$ shows advantage over classical. Note the scaling factor of $31e-$ 3 in the panel (d) for localization probabilities.
In our setup the localization phenomena is expected to occur after certain amount of time steps at a single defect node on a finite 2D lattice of size $\sqrt{N}\times\sqrt{N}$ unit cells starting from initial state $|i\rangle$. We apply the periodic boundary condition in order to avoid the appearance of edge states. The defect node is marked by parameters $\Theta_{1,2}^{def}$ that are different from surrounding $\Theta_{1,2}$. To estimate the possibility of given quantum walker to localize efficiently with high probability, we firstly scan the entire parameter range $\Theta_{1,2}$ for a given defect values $\Theta_{1,2}^{def}$. The measure of localization probability that is relevant for quantum search applications is defined as a peak $\max_{t\in[0,T]}P_{def}(t)$ in time-dependent density of the system state at the defect node,
@@
where $|d\rangle=\sum_{\sigma=\uparrow,\downarrow}|x=def,\sigma\rangle$ is the defect coordinate operator with both components of spin $\sigma$, and $|i\rangle$ is the initial state of the system defined as $|i\rangle=\frac{1}{\sqrt{2N}}\sum_{\sigma,x}|x,\sigma\rangle$ with summation over entire system. The plot of the maximum value of density $P_{def}(t)$ on defect over the interval of time $T=1000$ for test system size $N=40\times 40$ sites is shown in Fig. 1 for $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$. In Fig. 2(a) we show the color map representation of the same localization probability distribution on the top of phase diagram in parameter space. For other positions of defect $\Theta_{1,2}^{def}$ in parameter space the colormaps of localization probability are shown in Supplemental material [23].
@@
The islands of efficient localization in parameter space fill a small areas placed in two distinct types of regions: inside topological phases and on phase separation lines. We compare the search time and localization probability behavior with growing system size for the walker parameters marked by ‘1’ and ‘2’ in the Fig.2(a), that realize maximal values of probability within its regions. In Fig. 2(b), the search time is plotted against the square root of the system for the walkers ’1’ and ’2’. With the growing system size the behavior of search time demonstrates qualitative change - from $T\sim\sqrt{N}$ dependence to $T\sim\textit{const.}$. for both walkers. As we noted previously, such such a dependence of localization maximum on time was observed only for the Dirac QRW with particular type of Coulomb-like disorder potential [13, 14]. In Fig. 2(c) the localization probability of walkers ‘1’ and ‘2’ is compared to average density at non-defect nodes. We find that the constant search time regime corresponds to the $P_{def}(t)=O(1/N)$ with a large numerical prefactor. For the algorithmic comparison it is important to note that this decrease of localization probability with system size $N$ is faster than for non-topological QRW discussed in Ref. [10] $P_{def}^{Grover}=O(1/log\,N)$. The localization probability dependence in the form $O(1/N)$ arises due to the finite support of the trapped states. When the system size grows beyond the support of the trapped state, the search time asymptotic get saturated and we observe a clear dependence in the localization probability on the system size, see Fig. 2(c).
For the fixed parameters of quantum walker ‘1’ we visualized the distribution of localization probability in the space of defect parameters $\Theta_{1,2}^{def}\in[-2\pi,2\pi]$. The corresponding discussion in Appendix A shows that the localization probability depends only on $\Theta_{1}^{def}$ parameter, and is constant along the $\Theta_{2}^{def}$. Thus, it is more informative to present a dependence on $\Theta_{1}^{def}$ as shown in Fig.3(a) to compare with overlap criteria introduced below.
[…]
III Contribution of trapped states to dynamical localization
@@
To describe the origin of dramatically different localization behaviors for different parameters of the topological QRW and to explain why the regions of efficient localization represent only a small fraction of the parameter space, we analyze the spectrum and localization properties of eigenvectors of a unitary operator of the system with one defect.
The probability evolution at a defect site at discrete times $t$ can be rewritten through the unitary operator of QRW,
@@
$P_{def}(t)=\bigg{|}\langle d|\sum_{n,\lambda=\pm}e^{i\lambda E_{n}t}|n,\lambda\rangle\langle n,\lambda|i\rangle\bigg{|}^{2}.$ (8)
@@
The saturation of search time to constant scaling after system size becomes large enough suggests that the properties of states that mainly contribute to the Eq.(8) do not change with system size. This feature directly corresponds to the definition of trapped states previously introduced in Refs. [6, 15] in the context of QRWs staying localized around its initial starting node over infinite evolution time. We investigate the contribution of different eigenstates of $\bar{U}$ to the evolution of probability, $P(t)$, in Fig. 3(b) by comparing the product of overlaps $|\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle|^{2}$. The panels in Fig. 3(b) correspond to different parameters of defect (listed above each panel) denoted by vertical dashed lines for walker `$1^{\prime}$ in Fig. 3(a).
@@
Figure 3: (a) Comparison of maximum value of probability on defect for a walker ‘1’ normalized by numerically achieved maximum on the interval from Fig.2 as a function of defect parameter $\Theta_{1}^{def}$ with the maximal value of squared sum of overlaps defined in Eq.(9) with $\mathcal{M}$ containing one and two states respectively. All quantities are normalized to scale (0,1) for the purpose of comparison. The summation over two trapped states gives very accurate approximation of search probability peak. (b) The squared product of overlaps plotted as a function of energy of each eigenstate of operator $\bar{U}$ for walker parameters ‘1’ and defect parameter $\Theta_{1}$ indicated on top of each panel. Trapped states differ from other states by significantly higher value of overlap.
@@
The scale of $x$-axis in each panel of Fig. 3(b) shows that defect parameters with higher localization probability also correspond to higher overlap product for several states - in this case two pairs of particle-hole symmetric states. To further verify this observation, we compare the maximal value of overlap product
$W_{\mathcal{M}}=\left|\sum_{n\in\mathcal{M},\lambda=\partial}\langle d|n,\lambda\rangle\langle n,\lambda|i\rangle\right|^{2}$ (9)
@@
of a single eigenstate of $\bar{U}$ with the normalized probability dependence $max[P_{def}(t)]$ as function of $\Theta_{1}^{def}$. The comparison is made by taking a single pair of states that maximize overlap with set of indices =$\mathcal{M}=\{n_{max}\}$, as well as two pairs of states =$\mathcal{M}=\{n_{max},n_{second\,max}\}$. The results presented in Fig. 3(a) indicate that while =$\mathcal{M}=\{n_{max}\}$ selection overestimates the region of good search, the =$\mathcal{M}=\{n_{max},n_{second\,max}\}$ very precisely describes the high search probability peak. To further check this correspondence between the evolution of described by two pairs of states with maximal overlap and the maximum probability distribution, we show the same comparison for a defect $\Theta_{1,2}^{def}=(5\pi/8,\pi/2)$ along phase separation line and the line with walker `$1^{\prime}$ in Appendix B. The correspondence for a test system size with $\sqrt{N}=40$ shows that efficient search for such split-step QRW is mostly determined by the appearance of two pairs of trapped states that have large overlap with both defect and initial state.
@@
where we took into account the particle-hole symmetry relation $w_{-,j}^{\sigma}=(w_{+,j}^{\sigma})*$. It is important to emphasize that after calculating probability at a defect site, one finds that the two spin components exactly cancel oscillating parts of each other for a single value of $j$. That happens due to the different phases of oscillations between $sin$ and $cos$ in Eqs.(10) and (11) above, and results in:
@@
with $\text{sign}(\sigma)=\pm$. Thus, the contribution of each single pair of trapped states is constant and due to the initial condition of uniformly distributed state over entire system it could not describe the high peak in localization probability.
Figure 4: The comparison of density evolution at searched defect node (blue solid line) and another distanced node without defect (gray dashed line) with time for a walker parameters ‘1’ and defect marked by $(5\pi/8,\pi/2)$. The red dash-dotted line corresponds to the evolution described by only two pairs of trapped states that maximize product of overlaps and given by Eq. (III). The evolution from trapped states approximates with very good precision the first few peaks in probability and the time of their occurrence.
@@
The period of this oscillation is different from the fast oscillations of each pair of trapped states and is defined by the energy difference between the positive energy levels of two pairs of trapped states $(E_{1}-E_{2})$. Fig. 4 shows that expression in Eq. (III) captures very well the structure of oscillations of localization probability despite its simplicity. In Fig. 4 the results are presented for a walker ‘1’. We show the equivalent result walker ‘2’ in Appendix B. To further verify the agreement of predicted period of oscillations with the actual oscillation of probability on a defect node we used numerical Fourier transform. The agreement with period defined by $(E_{1}-E_{2})$ has errors below $2\%$ on the timescale of evolution $T=5000$ for system with $\sqrt{N}=40$. It was observed [24, 10] that in non-topological QRWs, that realize localization mechanism similar to Grover’s algorithm, and, the period of oscillation is defined by an energy difference within the same particle-hole symmetric pair of states at lowest energies. The difference between two definitions of the period shows that the mechanism of localization in topological QRWs is different comparing to non-topological QRWs studied in literature.